Find the special solution vector S in the nullspace of the the matrix A below that corresponds to the first free column 2 -10 16 3 3 A6-30 48 6 12 4 -20 32 -3 18 S1 S2 S3 S4 S5-
Answer #1

The null space of A is set of solutions to the equation AS = 0.To solve this equation, we will reduce A to its RREF as under:

Multiply the 1st row by 1/2

Add -6 times the 1st row to the 2nd row

Add -4 times the 1st row to the 3rd row

Multiply the 2nd row by -1/3

Add 9 times the 2nd row to the 3rd row

Multiply the 3rd row by 1/3

Add 1 times the 3rd row to the 2nd row

Add -3/2 times the 3rd row to the 1st row

Add -3/2 times the 2nd row to the 1st row

Then the RREF of A is

1

-5

8

0

0

0

0

0

1

0

0

0

0

0

1

If S = (s1,s2,s3,s4,s5)T, then the equation AS = 0 is equivalent to s1-5s2+8s3 = 0 or, s1 =5s1-8s3, s4=0,and s5= 0. Thus, S = (5s2-8s3,s2,s3,0,0)T = s2(5,1,0,0,0)T+s3(-8,0,1,0,0)T. This is the special solution of the equation AS = 0.. Here s2 and s3 are arbitrary real numbers.